Circular No. 7759

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001gc IN UGC 3375 M. Schwartz, Cottage Grove, OR; and W. D. Li, University of Califor-

nia at Berkeley, on behalf of LOTOSS (cf. IAUC 7514), report the discovery of an apparent supernova on an unfiltered image taken with the 0.5-m Tenagra III automated telescope on Nov. 21.4 UT (mag ∼ 16.2). SN 2001gc is confirmed at mag ∼ 16.8 on an earlier image taken by the Katzman Automatic Imaging Telescope under poor conditions on Nov. 19.5. The new object is located at α = 5^{h}55^{m}26^{s}.14, δ = +51^{o}54′34′′.2 (equinox 2000.0), which is 7′′.3 east and 4′′.3 south of the nucleus of UGC 3375. A Tenagra III image taken on Nov. 14.4 showed nothing at this position (limiting mag ∼ 19.0).

SUPERNOVA 2001fz IN NGC 2280 T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2001fz (cf. IAUC 7753), obtained by K. Rines on Nov. 19.50 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-II supernova.

COMETS C/2001 U8, C/2001 U9, C/2001 U10, C/2001 V1, C/2001 V2

Further to IAUC 7750, D. Hammer reports his measurements of additional Kreutz sungrazing comets found on SOHO website images by R. Kracht (C/2001 U8, C/2001 U10), X. M. Zhou (C/2001 U9), T. Scarmato (C/2001 V1), and A. and M. Mimeev (C/2001 V2). All the objects were seen in the C2 coronagraph. According to D. Biesecker, C/2001 U9 was a bright comet, visible from the edge of the C3 field-of-view until it disappeared behind the C2 occulter; it appears to have saturated the detectors in both telescopes. The reduced astrometry and orbital elements by B. G. Marsden appear on the MPECs cited below.

Comet  2001 UT  α_{2000}  δ_{2000}  MPEC
C/2001 U8 Oct. 18.601  13^{h}25.^{m}9  −10^{o}37′  2001-V39
C/2001 U9  22.096  13 14.1  −13 47  2001-V39
C/2001 U10  31.021  14 15.6  −15 12  2001-V60
C/2001 V1 Nov. 1.413  14 23.8  −15 54  2001-W47
C/2001 V2  4.132  14 31.9  −16 31  2001-W47
2001 November 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green